On January 4, Year 1, Larsen Corp purchased 10,000 shares of Warner Corp for $119,000 plus a broker's fee of $2,000. Warner Corp has 50,000 common shares outstanding and it is presumed the Larsen Corp will have a significant influence over Warner Corp. During Year 1 and Year 2, Warner Corp declared and paid cash dividends of $0.85 per share. Warner Corp's profit was $72,000 and $67,000 for Year 1 and year 2, respectively. The January 12, Year 3 entry to record the sale of 5,000 shares of Warner Corp for $65,000 should be
Answer #1
Percentage of ownership 20% =10000/50000
Purchase cost of investment 121000 =119000+2000
Add: Share of net profit 27800 =(72000+67000)*20%
Less: Dividends received -17000 =10000*0.85*2
Carrying value of investment 131800
Carrying value of investment sold 65900 =131800*5000/10000
Loss on sale of investment 900 =65900-65000
The January 12, Year 3 entry to record the sale of 5,000 shares of Warner Corp for $65,000 should be:
Debit Cash $65,000 ; debit Loss on sale of investment $900 ; credit Long Term investments $65,900